OpenStudy (anonymous):

Given the geometric sequence where a1 = 4 and the common ratio is 3, what is the domain for n?

OpenStudy (rock_mit182):

\[a _{n} = a*r^{n-1}\]

OpenStudy (rock_mit182):

since a=4 (first term) and r=3 (the ratio)

OpenStudy (rock_mit182):

\[a_{n} = 4*3^{n-1}\]